There are no Open Network Issues Currently
Upgrading server (Resolved) Medium

Affecting Server -

  • 17/06/2021 18:08 - 19/06/2021 18:00
  • Last Updated 07/07/2021 12:03

We are upgrading the server. there should not be much downtime in this period. 

Server update (Resolved) High

Affecting Server -

  • 06/11/2020 19:00 - 07/11/2020 02:00
  • Last Updated 06/11/2020 19:50

Update 06.11.20 19.50: The work are now complete. 

Update 06.11.20. 19.00: We are now starting the work. 


We are going to update the server

We don't expect any issues regarding this update, but still we want setup this service window. 

Our plan is to stop the MySQL and Email services on the server in case things doesn't go as planed and we need to do a rollback. 

Email will be queued for delivery once the update is complete by the sender.

Websites using MySQL databases will not be working as the service is shutdown in this service window, static webpages will work as normal.

There may be several reboots on the server in this service window. 


If there is any issue please open a ticket and we will take a look at it as soon as possible. 

Uplink maintenance (Resolved) Low

Affecting Server -

  • 29/04/2020 05:30 - 29/04/2020 05:45
  • Last Updated 29/04/2020 12:15

The maintenance will require a short interruption of the Internet uplink, we expect an offtime of less than 15 minutes.

unstability on server (Resolved) Low

Affecting Server -

  • 12/02/2020 16:10 - 14/02/2020 10:40
  • Last Updated 14/02/2020 10:42

12.02.20 - 16.10: We are aware of some unstability on the server. we are investigating. 
rebooting server. 
16.55: some processess did not respond and the system generated a high load. after the reboot it seems that the system is at a normal state. 
17.30: Seems that the issue is still there. we are working on the issue.
17.45: found a problem and killed the processes. seems to be running stable for the last 45 minutes. we will continue watching the servers load.


19.00: back to normal.


13.02.20 - 14.00: Still some problems with the server, we are watching the server. 


17.30: Doing another reboot. 


14.02.20 - 10.40: Seems that the server is running without any problems. 

Upgrading the system (Resolved) High

Affecting Server -

  • 07/02/2020 11:09 - 07/02/2020 11:20
  • Last Updated 07/02/2020 13:59

rebooting server for system update. 

Kernel update (Resolved) Medium

Affecting Server -

  • 22/10/2019 13:53 - 22/10/2019 14:05
  • Last Updated 22/10/2019 14:03

reboot of server after kernel update. 5 min downtime. 

prod 03 and prod04. 

kernel update (Resolved) Low

Affecting Server -

  • 15/10/2019 13:21 - 15/10/2019 13:31
  • Last Updated 17/10/2019 09:40

Upgrading kernel on server. A quick reboot of the server.  

Kernel update (Resolved) Medium

Affecting Server -

  • 09/10/2019 13:25 - 09/10/2019 13:31
  • Last Updated 09/10/2019 13:30

reboot of server after kernel update. 5 min downtime. 

Kernel update (Resolved) High

Affecting Server -

  • 18/09/2019 21:36 - 18/09/2019 09:50
  • Last Updated 18/09/2019 22:03

Kernel update on the server. A quick reboot after the update is finished. 

Upgrading the system (Resolved) Low

Affecting System - Customer system

  • 06/09/2019 15:23 - 07/09/2019 11:00
  • Last Updated 07/09/2019 15:25

We updated the customer system and migrated the system over to a standalone server. 

system upgrade (Resolved) Medium

Affecting System - main website

  • 14/11/2018 12:30 - 14/11/2018 13:15
  • Last Updated 14/11/2018 13:15

we are upgrading the main website. there will be some downtime on the site. 

Will not affect customers webpages.

Knowledebase Articles. (cPanel) (Resolved) Low

Affecting Other - Knowledgebase

  • 31/05/2017 00:00 - 14/11/2018 11:37
  • Last Updated 14/11/2018 11:37

Knowledgebase articles are not up to date. we are working on getting them up to date. 
If you have any questions or issues please open up a ticket to support. we will try and help you as good as we can!

Open a ticket to support if you got any questions about cPanel and the new control panel.

reboot of server (Resolved) High

Affecting Server -

  • 11/05/2018 16:00 - 11/05/2018 16:30
  • Last Updated 11/05/2018 18:00

Upgrading server, must do a reboot. 10 minutes downtime.


Reboot of server (Resolved) Low

Affecting Server -

  • 15/09/2017 09:39 - 15/09/2017 09:59
  • Last Updated 15/09/2017 10:02

We will reboot the server for updates. expected downtime for all sites on that server is 10 minutes.

reboot of server. Will be down for 2 minutes. (Resolved) Low

Affecting Server -

  • 21/06/2017 09:52 - 21/06/2017 09:55
  • Last Updated 21/06/2017 09:55

we will reboot the server. downtime will not be any more than 2 minutes.


migrating to new server with cPanel (Resolved) Low
  • 20/05/2017 00:57 - 16/06/2017 13:42
  • Last Updated 16/06/2017 13:42

We are migrating to the change will be a new control panel. cPanel. 
When we are migrating you we will inform you on email / SMS. please allow a couple of hours to let the DNS replicate across the network. 

We are doing this manually. Please open up a support ticket if there is any problem. 
Updates will be posted under here: 

24.05.17: most of the sites are moved over to the new server. 
Please log in to your controlpanel to check that everything is as it should be. 

Passwords on the email accounts are reset and you will need to use your full email address as username. 


31.05.2017 - 21:27
All sites is migrated to the new server. please open a ticket if there is any issues. 
secondary DNS / mail and has been shut down. 
(downgrading priority from high to low)


16.06.2017: legacy servers are shut down. closing issue.

Mail delivery (Resolved) Low

Affecting Server -

  • 02/06/2017 00:14 - 11/06/2017 16:44
  • Last Updated 11/06/2017 16:44

Please open a ticket to support if the server rejects email that is not spam. 

we are tuning the spam filter and want your feedback if we block mail that is not spam.

Planned Maintenance on May, 31st 2017 (Resolved) Medium

Affecting System - Domain registration

  • 31/05/2017 19:00 - 31/05/2017 23:19
  • Last Updated 31/05/2017 23:20

on 31 May 2017, we will carry out system maintenance on our production systems starting from 7 p.m CEST to account for the implementation of various major & minor updates.

All the systems will be online exept domain info / registering. 

The maintenance should last a few hours.


The scheduled system maintenance has been completed.
As of this moment, all services are up & running

Config fault on server (Resolved) Critical
  • 21/03/2017 10:00 - 21/03/2017 10:23
  • Last Updated 21/03/2017 10:53

we updated the config on the server. when we did that all virtual servers went to the first configured website... 

we updated all the config and the issue should be resolved now.

Networking software upgrade (Resolved) Low
  • 16/03/2017 01:00 - 16/03/2017 09:00
  • Last Updated 16/03/2017 09:12

We will be performing an upgrade of networking software.

During the maintenance window the server will lose networking connectivity for a period of up to five minutes. We expect the actual disruption to be minimal.

server down - (Resolved) Medium
  • 27/02/2017 11:40 - 27/02/2017 11:59
  • Last Updated 27/02/2017 13:03

server down -

power issue. switches unresponsive. switch replaced.
still working on the issue.

hard reboot on the server to get it back up.
soft reboot to fix a couple of things.

Upgrading customer system (Resolved) Medium

Affecting System - costomer portal

  • 16/12/2016 09:40 - 16/12/2016 09:50
  • Last Updated 16/12/2016 10:21

we are upgrading the customer portal. website will be down in this period.

SSL certificate (Resolved) Low
  • 10/11/2016 10:58 - 12/11/2016 00:00
  • Last Updated 11/11/2016 10:08

we have changed the SSL certificate on server. 

you will be asked to approve a new certificate on imap TLS login and smtp login

mail delivery, mail ending up in spam (Resolved) Low
  • 26/09/2016 08:21 - 06/10/2016 16:46
  • Last Updated 06/10/2016 16:46

Mail sent from are ending up in spam folder at some providers. 
we have put in place a temporary solution for delivery, but if the mail it self are considered to be spam this will have no effect! 

Please contact support if there is any problems, from address and to address and the time the mail was sent.

Update 05.10.16:
it seems that the issue is somewhat resolved. please open tickets with header info of mails that are not getting delivered.

Update 06.10.16:
seems to be Microsoft Outlook's spam filter. there are no errors in the logs. all mail are relaying as ther should.

unable to provision new services (Resolved) High
  • 04/08/2016 17:15 - 10/08/2016 12:46
  • Last Updated 10/08/2016 12:46

The issue is resolved


we are unable to provision new virtual servers on our shared hosting platform.

Temprary removed the possibility to pay when ordering shared hosting plans in order to prevent failed creation on a new account
enabled payment, seems to provision for now. we are still looking in to the issue. 

hopfully everything will work again soon.
no services are affected.

The problem will be taken care of ad-hoc. we will maually pick up profisioning that fails.
There may be an issue with adding / deleting users, we will help with that until the problem is resolved. 

Open a support ticket if there is an issue.

Reboot of master prod01 server (Resolved) Medium
  • 03/05/2016 13:00 - 03/05/2016 13:05
  • Last Updated 03/05/2016 13:05

we are going to reboot our prod01 server.

Websites will be offline for a couple of minutes.

enable ipv6 on server (Resolved) Low
  • 10/02/2016 10:31 - 10/02/2016 10:41
  • Last Updated 10/02/2016 10:41

Enabled on server, DNS zones has been updated. 

Enable IPv6 on server.

enable ipv6 on server (Resolved) Medium
  • 10/02/2016 09:15 - 10/02/2016 11:15
  • Last Updated 10/02/2016 10:25

Enabled DNS zones too. Closing issue. 


Update: IPV6 is up and running on the server. Adding IPV6 DNS Zones

we are adding ipv6 support on the server. there will be downtime within this timeframe.

hotmail / (Resolved) Medium
  • 05/02/2016 16:21 - 12/02/2016 00:00
  • Last Updated 09/02/2016 10:04

Update 09.02.2016: Seems to be OK now, closing network issue. 

Update 06.02.2016: We have disabled the relay and are starting to deliver mail directly from the server. 
please update us with a support ticket if any mail to, are beeing returned. Please add the error message in the ticket. 

Our are blacklisted at outlook /

we are using our secondary server to relay mail in this period. 

Please open a ticket to if you have any problems.

warnings on SPF and some other DNS issues is expacted. This is not a solution but a temp workaround.

Rebooting server (Resolved) Medium

Affecting Other - servers: and

  • 09/02/2016 09:49 - 09/02/2016 10:15
  • Last Updated 09/02/2016 10:04

we are rebooting the secondary servers, and

will not affect systems. 

Migrating fr01 to prod01 (Resolved) Medium
  • 17/11/2015 19:49 - 22/11/2015 23:50
  • Last Updated 22/11/2015 23:52

We are in the prosess to migrate server to server.

we will run a backup and restore.

If the migration fails we can do a roll-back as nothing is deleted in the prosess. 

Open a ticket if there is a problem / question. or your customer portal.


Update: Migration is done, please login to your control panel to chack your DNS config.

Final update: everything seems to be in place. please open a ticket if you have any problems. - webserver halts at 21:25 (Resolved) Medium
  • 04/11/2015 10:29 - 20/11/2015 21:00
  • Last Updated 19/11/2015 22:54

webserver halts at 06:25 in the morning
one process is not killed and apache wont start. 
the process must manually be killed and restarted. 

we are investigating the issue and will continue to restart the service manually until the problem is resolved.

logrotate is causing the issue but the real problem are apache 2.4.10. 
we have moved the cron for logrotate to 21.25 at night. and will handle the alerts than.
until a fix in the stable release channel is available we will be handling this manually.
This also happens when a customer updates ther server in webmin / virtualmin (Only the Apache2 / PHP settings).

An alarm goes of after 3 minutes and we will most likely be online again after 10 minutes. 

DNS / FTP / Mail - is not affected.


Final update. FR01.EU server is retired. migrated to

Server downtime - Re-insatll (Resolved) Medium
  • 30/10/2015 19:38 - 01/11/2015 23:00
  • Last Updated 01/11/2015 22:27

The problem with the main server are too severe, we have to reinstall. This is a good thing tho, a fresh install will resolve the small bugs :)

This will affect all the virtual servers assigned to this server.

Email and DNS will still be served by our backup servers ( and but email accounts / webmail won't be accsessable.

we will first run a complete backup of the server. transfer it to our FTP backup server. 
when this is done we will shutdown mail server and web server including SQL servers. 
a incremental backup will be created and transfered to our FTP backup site. 

we will than begin the reinstall of the server.

For support please open a ticket (NOT email) to our support department. 

Progress and updates in this network issue:
GMT + (Oslo)
30.10.15 - 20:00: We begin the first complete backup of all the virtual servers assigned to fr01. 
31.10.15 - 00:00: We are continuing 31.10.15 at 21.00 to avoid peak time. we will start again then. - All systems are working as normal untill we restart the prosess.
31.10.15 - 12:30: Backups are running, it seems that we can push this critical work forward. We are now running at the last stage of a pre-reinstall. a few small things to do and an incremental backup and the reinstall begins.
31.10.15 - 21:00: Starting incremental backup. now. a new update when reinstall are beginning. Mail server and Web server server are shutdown now.
31.10.15 - 23:30: We are starting the reinstall now.
01.11.15 - 00:05: The server has a new install, we're beginning to restore systems.
01.11.15 - 01:35: we are beginning to restore servers from backup.
01.11.15 - 04:40: is back online. a few minor glitches. this will be taken care of in the next couple of days. downgrading severity from critical to medium.
01.11.15 - 21:00: Problems with DNS issues, not transfering DNS zones. we are working on it.

Last update: 01.11.15 - 22:20: Everything seems to be working fine. please create a ticket if you have any issues.

web server issues (Resolved) Critical
  • 28/10/2015 13:33 - 01/11/2015 02:00
  • Last Updated 01/11/2015 04:42

we have problems with apache on this server. it suddenly failes and does not restart the services. 
we are working on this issue. 

this issue is only on the web server part. Mail and DNS are working as before. 

GMT +1:
29.10 - 13:35: Rebooting server
29.10 - 16:30: Web server is stable, we are monitoring the server closly. the issue will be updated.
30.10 - 08:30: web server was unavailable again. again, mail / dns is not affected.

30:10 - 19:40: Please se the reinstall issue. the progress will be updated there.

Upgrading (Resolved) Medium

Affecting System -

  • 14/10/2015 18:00 - 14/10/2015 18:21
  • Last Updated 14/10/2015 18:00

We are upgrading the main site, most likly a downtime on the page for about 10 minutes.

Maintenance of secondary server (Resolved) Medium
  • 10/10/2015 20:00 - 11/10/2015 04:00
  • Last Updated 12/10/2015 09:17

The maintenance  will not have any impackt on services. it will be downtime on this server for about 20 minutes in this window. 

all services will continue as normal. main server ( and secondary #2 ( will still be online to serve DNS request and receve email. 

please open a ticket if you experience any issues after the maintenance.

Maintenance of secondary server (Resolved) Medium
  • 15/10/2015 04:00 - 15/10/2015 23:15
  • Last Updated 08/10/2015 09:36

The maintenance  will not have any impackt on services. it will be downtime on this server for about 20 minutes in this window. 

all services will continue as normal. main server ( and secondary #2 ( will still be online to serve DNS request and receve email. 

please open a ticket if you experience any issues after the maintenance.

loss of connectivity (Resolved) Critical
  • 01/10/2015 15:40 - 01/10/2015 15:45
  • Last Updated 01/10/2015 15:45

We lost connectivity in 5 minutes. everything is up and running again.

DNS servers (Slave) (Resolved) High

Affecting System - Slave DNS servers

  • 26/08/2015 10:33 - 26/08/2015 22:53
  • Last Updated 26/08/2015 22:53

We are refreshing all slave dns servers.
main dns server are still up and running, no config will change there. on the slave servers we will rebuild the database.
This will not have any impact on your services.

update: Working on
Update, there will take some time, all DNS requests are served by and so there will be no impact on services. 
Update: had to reboot and main server fr01

Update: the issue is solved. resyncing zones across the network. 
update: syncing completed. keeping this open to see if the transfer are syncing as normal ( one times every hour )

Final update: confirmed, DNS is back to normal. all servers are replicating DNS zones as normal. 

currently blocket with / hotmail. (Resolved) Low
  • 13/08/2015 13:11 - 15/08/2015 05:00
  • Last Updated 17/08/2015 16:16

We are currently blocked by / / due to user reports.

We are looking in to it now. 

A temp end time is set, this will be updated when we know more. 


Update: we have contacted microsoft for removal.


Update: removal can take 24 - 48 hours to be completed. Setting priority to low.

lost connectivity to server (Resolved) Critical
  • 14/08/2015 01:58 - 14/08/2015 02:25
  • Last Updated 14/08/2015 02:29

we lost connectivity to

this is a backup server, there are two other servers that cover this server.

impact to services are none. 

The issue is resolved. 29 minutes downtime. 

Email to domains, (Resolved) Low
  • 27/05/2015 11:55 - 27/05/2015 22:25
  • Last Updated 29/05/2015 10:26

Problem with sending email to domains with /
we have contacted microsoft to have our servers ip removed. 
Updates will follow as soon as we know more.

Update: Microsoft has unblocked our mail server. please open a abuse ticket if there still is a problem.
this can take up ti 48 h to replicate to all of microsofts servers. 

Update: Confirmed resolved.

Slow loading on (Resolved) High

Affecting System -

  • 26/05/2015 07:18 - 26/05/2015 10:32
  • Last Updated 26/05/2015 10:32

The loading time on are about 20 seconds.

other sites are running as normal. 

we don't know why, but we are investigating the problem.


Problem fixed

unavailable DNS management server. (Resolved) Low
  • 27/04/2015 13:00 - 27/04/2015 19:06
  • Last Updated 27/04/2015 13:37 is a pure DNS management server. 

This will have no impackt on services. DNS on fr01, nl02 and will still be online and serve DNS queryes.

Expected to be online again 16.00 (GMT +1)

Lost connection to FR01 (Resolved) Critical
  • 13/04/2015 10:00 - 13/04/2015 10:14
  • Last Updated 13/04/2015 13:16

We lost connection to the server had to be rebooted. 

This issue is now resolved.

Upgrade of server (Resolved) Low
  • 17/04/2015 04:30 - 17/04/2015 11:20
  • Last Updated 10/04/2015 09:43

We are upgrading the secondary mail and DNS server. it will be down for a couple of minutes in this period.
This will have no impact of services as and server are still up.

moving DNS01 to another node (Resolved) Low
  • 31/10/2014 14:35 - 04/11/2014 11:31
  • Last Updated 04/11/2014 11:31

we are moving to another node. will be some downtime on

Shutdown of (Resolved) Low
  • 17/09/2014 19:16 - 30/09/2014 00:00
  • Last Updated 01/10/2014 10:50

We are shutting down
all services are moved to and servers. 

we start by shutting down DNS and mail servers, please open a support ticket if this affects you or your site.

The server is completly shut down.

Slave DNS/Mail server change (Resolved) Low
  • 10/09/2014 16:00 - 12/09/2014 00:00
  • Last Updated 12/09/2014 15:50

We are moving our Slave DNS/mail server to new server.

both servers are online and will continue to serve DNS requests.

in case of dns resolve/ update issues please open a ticket. 

Slave mail server dows not accept mail (Resolved) Low
  • 10/09/2014 20:00 - 11/09/2014 19:00
  • Last Updated 11/09/2014 10:14

Seems that the newly setup slave mailserver wont receive mail.
as for now this is not a problem.

we will continue to work on the issue

Server issue, downtime (Resolved) Critical
  • 02/08/2014 04:31 - 02/08/2014 10:07
  • Last Updated 02/08/2014 11:00

There was an error on the server witch was resolved. The autosystem did not reboot the server when the error was detected. 

The issue are now resolved 

power failure (Resolved) Low
  • 07/05/2014 11:00 - 07/05/2014 13:30
  • Last Updated 07/05/2014 11:32

Power failure on location where the DNS management server are located. 

Does not affect DNS. only management.

Update: Resolved. all services up and running. 

Email - google are blocking our ip. (Resolved) Medium
  • 11/04/2014 00:00 - 12/04/2014 10:00
  • Last Updated 12/04/2014 12:02

Google are blocking our mail server.
This is a temporary problem and will resolve itself when the traffic returns to normal.

goes through but it takes a little longer than normal

Update : problem should be resolved now.

Migrating server DE01 to FR01 (Resolved) High
  • 21/03/2014 20:00 - 21/03/2014 23:41
  • Last Updated 21/03/2014 23:41

Update 23:45: Migration are done. we have to wait for DNS replication.

Open a ticket if there is a problem. 


Migration from de01 to fr01.

we will shutdown the mail server on de01 and the secondary server will take over and store the email.

We hope this upgrade will go smoothly but will keep updating on the progress.

When we start the migration the main site will be moved to amazon with a static page for information. When the migration are done and DNS are updated the site will automaticly become available again. 

You can test to see if the server and DNS are updated by pinging your domain. 
if you see the old ip address the DNS are not updated, Yet.

Old IP :
New ip:

Update: 20.15 GMT +1
The migrate has started

Update: 23.00 GMT +1

Migration are complete. DNS records are beeing updated now. after that the replication of DNS zones will take a while. 

Domain registering and domain info (Resolved) High

Affecting System - Domain

  • 22/12/2013 14:35 - 23/12/2013 00:00
  • Last Updated 23/12/2013 00:33

23.12.13: 00:00 GMT +1. 
Our register came online again. the issue is resolved. 

22.12.13: 17:10 GMT +1. 
Still no updates. seems that our registrar are completly down. we are hoping they are online again very soon.

22.12.13: 14:30 GMT +1.

Seems that there are some problems with domain registering. our registars API are down. 

We are looking in to the problem.

This issue will be updated.

Affects TLD's:


Updating website (Resolved) Low

Affecting System -

  • 08/10/2013 13:30 - 08/10/2013 13:44
  • Last Updated 08/10/2013 13:45

We will update the main webpage 
should not be more than 10 minutes downtime on main page. all servers are up and running.

will not affect any customers. 
Client Portal will be unavailable while upgradeing.